An airplane accelerates down a runway at 3.20 m/s2 for 32.8 s until is finally lifts off the ground. Determine the distance traveled before takeoff.

1
Expert's answer
2022-03-18T09:15:52-0400

The distance traveled by airplane before it takeoff


"d=v_0t+at^2/2\\
=0\times 32.8+3.20\times 32.8^2/2=1720\:\rm m"